### v1

# New theoretical Fe II templates for bright quasars

Abstract
We present a set of new theoretical Fe II templates for bright quasars covering a wavelength range of 1000-10000 Å, based on the recent atomic database available in the C23.00 version of the photoionization code CLOUDY. We compute a grid of models for a range of incident photon flux and gas density and for multiple microturbulence velocities. We analyze the ratios of Fe II emission over a variety of wavebands and compare them with observations. Our key results are: (1) Despite the use of the newest atomic data we still confirm the long-standing problem that the predicted Fe II UV/optical ratio is significantly larger than that observed in the AGN spectra. (2) The ratio is not significantly affected by the variations in the microturbulence and the metallicity. (3) The turbulence can create an additional apparent velocity shift of up to 1000 km/s in the spectra. (4) There is no single Fe II template that can fit the observational data covering UV to optical wavelength range. We shortly discuss the most likely effects responsible for the Fe II UV/optical mismatch problem: the assumption of the constant column density and the assumption of the isotropic emission implying equal contribution of the bright irradiated faces and the dark shielded faces of the clouds.
1 Introduction
The dominant component of the bright active galactic nuclei (AGN) in the optical/UV band is the continuum component coming from the cold optically thick geometrically thin disk (see e.g. Krolik, 1999; Karas et al., 2021; Zajaček et al., 2023a, for reviews). Locally this component is well described as a power law although at high quality and/or in broadband data some curvature of the component is visible (e.g. Capellupo et al., 2015). Superimposed on this continuum component are numerous strong and broad emission lines like H, Mg II, C IV, and Ly , depending on the redshift range covered by the data. These lines are easy to identify in the individual spectra of AGN and quasar composite spectra (e.g. Francis et al., 1991; Vanden Berk et al., 2001; Telfer et al., 2002; Scott et al., 2004; Ren et al., 2022). These lines come from the Broad Line Region (BLR) which is located at a fraction of a parsec from the central supermassive black hole (SMBH). This region reprocesses the bulk of the radiation from the central parts very close to the SMBH.
Apart from these well-resolved strong individual lines, we have a contribution from transitions in weakly ionized heavy elements, predominantly iron. These transitions are so numerous that they form a pseudo-continuum well visible in the near-IR (NIR), optical and UV bands. This pseudo-continuum affects considerably the fitting of the strong individual emission lines from BLR. The issue has been identified a long time ago (Sargent, 1968; Collin-Souffrin et al., 1979; Wills et al., 1985, 1980). The important role of Fe II diagnostics has been recognized by Boroson & Green (1992) in their studies of the many AGN parameters using the Principal Component Analysis (PCA) approach. Their study led to the identification of the key role of the Eigenvector 1 (EV1) in the quasar sequence. The ratio, , of the equivalent width (EW) of the Fe II in the 4434-4684 Å wavelength range to the EW of H played a key role in EV1, and later it became the foundation of the optical Quasar Main Sequence plots showing the Full Width Half Maximum (FWHM) of H vs (Sulentic et al., 2000; Shen & Ho, 2014; Marziani et al., 2018; Panda et al., 2018, 2019b).
Since Fe II transitions are so numerous, certain standard templates were developed separately for the optical and for the UV bands which were used for detailed modeling of AGN spectra. Some of the templates were purely observational, i.e. created by selecting the object with relatively narrow lines and subtracting the known lines and continuum. The residual was considered as coming from Fe II (e.g. Vestergaard & Wilkes, 2001; Marziani et al., 2003; Park et al., 2022). Other templates were mostly theoretical, based on computing the transition rates (e.g. Bruhweiler & Verner, 2008; Kovačević-Dojčinović & Popović, 2015). Some templates were a combination of both approaches (Tsuzuki et al. 2006, but also partially Kovačević-Dojčinović & Popović 2015).
The main issue with the empirical or observational templates is that these templates are limited by the signal-to-noise ratio (SNR) and the wavelength coverage of the spectra. There is not yet a single empirical template that can cover the entire optical to UV bands simultaneously.
In this paper, we use the latest version C23.00 of the photoionization code CLOUDY (Chatzikos et al., 2023), which contains a large database of the Fe II transitions, to develop a new theoretical Fe II template that covers a wide wavelength range of 1000-10000 Å. This new template will be useful to fit the AGN spectra simultaneously from optical to UV bands. Some properties of Fe II emission, like the ratio of the Fe II to Mg II line in UV and the trends with the EV1 were studied by Sarkar et al. (2021); Dias dos Santos et al. (2024) but here we study systematically several correlations which can also be tested against observational data in the NIR to UV wavebands.
2 Method
We use the latest version C23.00 of the photoionization code CLOUDY (Chatzikos et al., 2023). CLOUDY code has four different Fe II atomic data sets reported by Verner et al. (1999), Bautista et al. (2015), Tayal & Zatsarinny (2018) and Smyth et al. (2019). A comparison of these four data sets is presented in Sarkar et al. (2021). For our model calculations, we use the Fe II data set of Smyth et al. (2019) which is the default Fe+ dataset in CLOUDY C23.00. The Smyth et al. (2019) Fe II data set includes 716 levels extended up to 26.4 eV generating 255974 emission lines.
We computed several grids of photoionization models that span five decades in the hydrogen gas density, 9 14, and the incident hydrogen-ionizing photon flux, 17 22, with a step size of 0.25 dex. We adopted a fixed hydrogen column density of 10 cm and set the abundance to the default solar abundance, as in Sarkar et al. (2021). For the description of the spectral shape of the incident radiation, we used the standard AGN.SED available in the CLOUDY database, which is similar to the continuum of Mathews & Ferland (1987).
We generated such models with different values of microturbulence velocity (V). We used six values of microturbulence velocity: 0, 10, 20, 30, 50, and 100 km/s. Such range for the microturbulence velocity was inspired by previous works (Baldwin et al., 2004; Bruhweiler & Verner, 2008; Panda et al., 2018). For a certain set of values of and , we did not get any Fe II emission. We listed these values in Table 1.
Having the set of computed spectra, we constructed several ratios of Fe II emission by integrating the spectra over the selected wavelength ranges. Such ratios are frequently studied in the observed spectra, and we aim to analyze the trends in these ratios as functions of the density and the ionizing flux.
We also examine the effect of metallicity on these ratios by varying metallicity as 2, 5, 7, 10 and 20 times the solar metallicity.
21.25 | 9.00 |
21.50 | 9.00 |
21.50 | 9.25 |
21.75 | 9.00 |
21.75 | 9.25 |
21.75 | 9.50 |
22.00 | 9.00 |
22.00 | 9.25 |
22.00 | 9.50 |
22.00 | 9.75 |
Note: This set of parameters is valid for all the turbulence velocities except for zero. In case of no turbulence velocity in addition to the upper parameters, Fe II emission was also not obtained for = 22 and = 10.
2.1 Definitions of optical and UV Fe II
Since the Fe II emission is of a broad band character the integrated value depends critically on the adopted range of integration. In our computations, we use several such ranges that were frequently considered in the literature.
To characterize the broadband emission in the optical range we integrate Fe II from 3500 to 7200 Å (Véron-Cetty et al., 2004). Some of the studies concentrate on much narrower ranges. For example, the range from 4434 - 4684 Å was introduced to parameterize the Fe II emission close and bluewards of the H line and to characterize the relative importance of the Fe II to H emission by the corresponding ratio (Boroson & Green, 1992). This ratio was well studied observationally in many papers (e.g. Osterbrock, 1977; Bergeron & Kunth, 1984; Shen & Ho, 2014; Śniegowska et al., 2018; Martínez-Aldama et al., 2021; Marziani et al., 2022).
Broad UV Fe II range can be well characterized by the Fe II emission integrated in the wavelength range from 1250 to 3090 Å (Vestergaard & Wilkes, 2001). This range contains the Fe II bump at 2500 Å seen in the spectra as well as in early models of Fe II emission (Wills et al., 1985).
We also conduct a systematic analysis of two UV and two optical wavelength ranges that may be used to determine the spectral shape. The UV ranges, 2650-2800 Å and 2800-3090 Å characterize the Fe II emission on each side of the Mg II line. The optical ranges are 4434-4684 Å and 5150-5350 Å corresponding to the Fe II emission at the blue and red sides of the H line, respectively. The Fe II emission on both sides of Mg II and H forms clear local bumps or peaks. Therefore, for convenience, we refer to these regions as blue and red wings (e.g. Sarkar et al., 2021).
3 Results
3.1 UV to optical Fe II emission
We calculated the ratio of the UV to optical broadband Fe II emission with the integration range specified in Section 2.1. The results are shown in Figure 1 using contour plots. The representative value in the central parts of the plot is 0.7 in the log space, i.e. the ratio is typically of the order of 5 in linear space. As already argued by Joly (1981) and Wills et al. (1985), a ratio of the order of 1 in the relative intensities of the optical to UV lines implies a column density of about to cm to balance Fe II line optical depths and the Balmer opacity. This value is only weakly sensitive to the ionization flux or density. However, we see that this ratio rises rapidly when either the ionization parameter becomes very high or very low, and the cloud density rises. At higher values of the ionization parameter, above cm s, this steep rising wall forms a linear trend in the diagram, with the characteristic position correlated with the density. The ratio at this wall exceeds 100. At low values of the ionization flux, below cm s, the ratio depends mostly on the density, changing from the factor 2.5 in the linear scale for densities below cm to values above 100 for densities above cm. Lines are almost vertical there, so for low values of the ionization flux this ratio can serve as a density indicator. At intermediate values of the ionization flux, above cm s, the dependence on the two parameters is more complex. Thus, when we do not have an estimate of the ionization parameter in the source, high values of the UV to optical Fe II emission can indicate either very high or very low ionization flux. Sarkar et al. (2021) stressed the role of the turbulent velocity in the Fe II intensity but the global ratio of the UV to optical is relatively unaffected.
The ratio shown in Figure 1 represents the expectations based on the assumption of a single zone responsible for both UV and optical emission, which may not be the case in actual objects. We address this issue in the discussion.
3.2 Red-to-blue UV Fe II wings ratio
In our next plot, we concentrate on the global properties of the UV part of Fe II emission which is important for the proper modelling of the Mg II line. The Fe II emission under the Mg II line is never close to zero, even right under the central parts of the Mg II, at 2800 Å, as adopted in the creation of UV Fe II observational templates (e.g. Vestergaard & Wilkes, 2001). However, due to the overlap of the Fe II and Mg II there, it is much easier to measure the strong wings that Fe II forms at both the red and blue sides of Mg II. Proper modeling of the underlying Fe II may affect the conclusions about the dynamics of the Mg II emitting region (e.g. Jonić et al., 2016; Popović et al., 2019).
The contour plots for the ratio of red (2800-3090 Å) to blue (2650-2800 Å) wings of UV Fe II are shown in Figure 2. The plot of the ratio shows a complicated structure. We again see a trend similar to the one displayed in Figure 1. The same values of the ratio are found at low-density high ionization flux and high-density low ionization flux. At low ionization flux, the ratio again depends mostly on the density. However, in this case, we additionally observe the dependence on the turbulent velocity, as in the central parts the ratio is initially rising, and finally decreasing when the turbulent velocity crosses the value 50 km s. Maybe this is connected with the fact that the overall range in the measured ratio is much smaller for the red to the blue wing (typical range between 0.3 and 1.8) than in the case of the entire UV to optical range covering a broad range of magnitudes.
3.3 The effect of the UV Fe II broadening on the measurements of the outflow properties
In the analysis of their Sloan Digital Sky Survey (SDSS) sample, Kovačević-Dojčinović & Popović (2015) noticed a systematic averaged redshift in the UV Fe II to the optical Fe II. To check if such a shift can be caused just by the shape of the Fe II emission when broadened with the turbulence effect included, we used a specific Fe II template characterized by the ionization flux, , and the density, . We applied the Gaussian smearing to the template, and we determined the peak wavelength of the blue wing of UV Fe II. The peak is usually located close to Å , and it is dominated by the multiplets 62 and 63 (see Kovačević-Dojčinović & Popović 2015). However, when we model the entire Fe II emission of the blue wing we see that the position of the peaks moves with the adopted smear velocity. We plot this effect in Figure 3 simply in Å units, and in Figure 4 in velocity space. The shift corresponds to the apparent velocity from 100 km s up to 300 km s when the turbulence is not included, and the presence of the turbulence affects the apparent direction of the outflow and its amplitude. Apparent velocity shifts up to 1000 km s are possible even though no velocity is present in the model which corresponds in each case to the rest frame spectrum. In Kovačević-Dojčinović & Popović (2015) more than half of the sources show a kinematic offset above 1000 km s. We do not see such large shifts due to the Fe II smoothing but we experimented only with a single template.
3.4 Red-to-blue optical Fe II wings ratio
This part of Fe II emission is important for accurate modelling of the H zone, as recently stressed by Popović et al. (2023). Proper modelling of the Fe II pseudo-continuum may strongly affect the conclusions about the dynamics of the BLR (like the presence or absence of the very broad line component to H, and the role of the intermediate line region), and the modelling of the quasar main sequence (Panda et al., 2018, 2019a, 2019b; Marziani et al., 2021).
We show contour plots for the ratio of red (5150-5350 Å) to blue (4434-4684 Å) parts of optical Fe II in Figure 5. This ratio shows again the non-monotonic dependence of the ionization parameter but interestingly, almost no dependence on the density. The minimum value of around 1.0 is present at the ionization parameter cm s, increasing when we move both up and down in the ionization flux. The dependence on the turbulent velocity is marginally noticeable, particularly for this minimum value. Some traces of dependence on the density are only seen for very low ionization flux, below cm s, and only for very high densities. The value of the turbulent velocity affects the extension of this region.
Abundance (in ) | , | UV Fe II/optical Fe II | UV red/blue wings | Optical red/blue wings | R |
1 | 18,10 | 2.56 | 2.13 | 1.81 | 0.91 |
20.5,12 | 8.98 | 1.26 | 0.84 | 1.09 | |
5 | 18,10 | 2.28 | 2.05 | 1.99 | 1.27 |
20.5,12 | 5.63 | 1.21 | 0.75 | 2.49 | |
20 | 18,10 | 3.03 | 1.77 | 1.88 | 0.99 |
20.5,12 | 4.94 | 1.03 | 0.76 | 4.42 |
3.5 in the optical band
When computing the Fe II emission with CLOUDY we also have, as a by-product, the intensities of other emission lines. The ratio of the optical Fe II emission to H plays an exceptional role in the EV1 of the quasar sequence (Boroson & Green, 1992; Sulentic et al., 2000; Shen & Ho, 2014; Marziani et al., 2018; Panda et al., 2018, 2019b; Du & Wang, 2019; Panda & Marziani, 2023). We estimated as the ratio of the integrated optical Fe II flux in the standard wavelength range 4434-4684 Å to the flux of H at 4861.32 Å. We present the corresponding contour maps in Figure 6.
These plots show that most of the map area corresponds to the value of smaller than 1. Higher values, above 3, are only predicted for the intermediate values of the turbulent velocity (20 - 30 km s, and only for very high values of the ionization parameter () which supports the earlier findings by Panda et al. (2018) and Panda (2021). Smaller values of correspond to the intermediate A-type quasars. The basic classification of quasars into class A and B, introduced by Sulentic et al. (2000) was analogous to the older classification of Seyfert galaxies into Narrow Line Seyfert 1 (NLS1) class and other Seyfert 1 galaxies (Osterbrock & Pogge, 1985), only the limiting velocity was different (4000 km s vs. 2000 km s), which reflected the systematic difference in the mass of the central black hole. In both cases, one class (NLS1 and type A) of objects corresponded to relatively higher Eddington rate sources than the other class, which was reflected in line properties (intensities and shapes). Later, the specific ranges of were used to explore full 2-D classification (see e.g., Marziani et al., 2018). A-type quasars with were named Extreme Population A (xA) sources, and they were found to be applicable as standard candles for cosmology (Marziani & Sulentic, 2014; Dultzin et al., 2020; Marziani, 2023; Panda & Marziani, 2023). Most of the xA sources have the values of between 1 and 2 (Marziani et al., 2018) but Śniegowska et al. (2018) found one source with a large ratio. A few more such sources were reported by Shen et al. (2011) but as discussed in Śniegowska et al. (2018), those measurements were not of high quality. Moreover, the sources with 2 are reported in the newer catalogs (e.g. Negrete et al., 2018; Rakshit et al., 2020, 2021; Wu & Shen, 2022).
We, thus, could conclude, from Figure 5, that xA sources require high-density BLR () and very high radiation flux. However, our templates were composed of solar metallicity. Some authors argued (e.g. Panda et al., 2019b; Śniegowska et al., 2021; Panda, 2021; Marziani et al., 2024) that these sources, in addition, require highly super-solar metallicity.
We compared these predictions with the statistical distribution of the reported on the basis of SDSS quasar studies. We used two data sets: older massive fits to the SDSS quasars, based on data release 7 (DR 7), were reported by Shen et al. (2011). Newer results, based on DR 14, were reported by Rakshit et al. (2020). The results are based on the spectral decomposition of the Fe II emission integrated within the traditional range 4434-4684 Å (Boroson & Green, 1992). In both cases, we selected only the sources with reliable spectra decomposition by requiring that the relative error of the FWHM of H and the measurement are below 20%. The resulting histograms, shown in Figure 7, consistently show the peak of the values at about 0.6 in both samples. There are very few objects with values above 4.0 which we ignored for better clarity of the plot. Such values are present either in the region of higher density and high ionization flux, above cm s, or low ionization flux and low density (lower left corner of plots in the Figure 6). The first location seems more likely as it is smoothly connecting to even larger values of than the average value, and the observational histogram also smoothly extends to higher values. However, we should stress that the observational results are not based on the Fe II emission models presented in this paper. Results of both Shen et al. (2011) and Rakshit et al. (2020) are based on the use of the optical Fe II template from Boroson & Green (1992).
3.6 The effect of metallicity
To investigate the effect of metallicity on Fe II emission, we rebuild our models for abundances equal to 2, 5, 7, 10, and 20 times solar abundances, with a fixed microturbulence velocity of 20 km/s. We plot contours for the ratios estimated in the previous sections in Figures 8-11 for various abundances. The values of these ratios at two arbitrary positions are given in Table 2 to quantify their variations with abundances.
The UV Fe II to optical Fe II contours, shown in Figure 8, exhibit marginal changes with abundances. The ratio decreases with increasing abundance at (=(20.5,12), however at no such trend is seen at lower incident flux, and lower density (=(18,10). Verner et al. (2003) investigated the effect of varying abundance on the ratio of UV Fe II to optical Fe II and found that the ratio decreases from 8 to 5.6 when abundance increases from solar to 10 times solar abundance for and a microturbulence velocity of 1 km/s. They concluded that the optical Fe II emission increases more than the UV Fe II emission when the abundance increases which is supported by the findings from Panda et al. (2019a).
Figure 9 shows the contours for the ratio of UV wings, whereas Figure 10 shows those of optical wings. As seen from the Figures and Table 2, both ratios are essentially unaffected by the abundance.
The contours for R are plotted in Figure 11 which shows small variations with abundance. The ratio increases with increasing abundance at high ionization flux and high density (=(20.5,12), but no such trend is seen at (=(18,10).
3.7 Fe II templates
For convenient use in future spectral fitting, we created a new family of Fe II templates. They come from the simulations as described in Section 2, i.e. those are 2646 models covering the range of the local density, ionization flux and turbulent velocity. Output files from CLOUDY are usually large, and data fitting does not require full resolution available in these output files since most of AGN have broad Fe components. Therefore, we binned the output spectra using the 2 Å bins. We tested whether this does not lead to a considerable problem in data fitting by assuming that the AGN spectrum comes from the medium with the velocity dispersion 2000 km s and we compared the Fe II spectrum in the 2650 - 2850 Å range without any prior binning, and with a bin size of 1, 2 and 5 Å. The result is shown in Figure 12. Up to the bin size of 2 Å there is no net difference in the broadened spectra but the bin size of 5 Å seems to imply already a systematic difference, so we adopted the 2 Å bin size optimizing between the accuracy and the size of the file. The complete set of Fe II templates for solar metallicity is available through zenodo111https://zenodo.org/doi/10.5281/zenodo.10532690 and GitHub222https://github.com/Ashwani-88/Fe2_template links.
Source | RA | Dec | Redshift | log(L[erg/s]) | log ( [M]) | Reference | |
HE 0413-4031 | 1.39117 | 46.74 | 8.87 | 1.66 | Prince et al. (2023) | ||
SDSS RM102 | 0.86114 | 45.00 | 7.92 | 0.329 | Shen et al. (2019) |
The redshift to this source was recently firmly established from the narrow [OIII] line using the near-IR spectrum (Zajaček et al., 2023b).
3.8 Exemplary cases
We test how the new Fe II templates compare with the observed AGN spectra. Since we use the SED more appropriate for bright quasars to construct the Fe II templates, we selected two exemplary sources of this type namely, HE 0413-4031 and SDSS RM102. The basic properties of these sources are mentioned in Table 3. In the case of HE 0413-4031, only narrow UV band data are available while the spectrum of SDSS RM102 continuously covers the broad UV-to-optical band.
3.8.1 HE 0413-4031
This quasar has been observed with the Southern African Large Telescope (SALT) for 13 years with the aim of performing the reverberation mapping of the Mg II line and the UV Fe II (Zajaček et al., 2020; Prince et al., 2023; Zajaček et al., 2023b). The obtained spectra cover a narrow range around the Mg II line but the spectral shape did not show any strong variability apart from the normalization of the spectrum. We thus constructed the mean spectrum of this quasar by combining all 32 available data sets for a better S/N ratio. The shape of the Mg II line in this source is particularly simple, well fitted as a single kinematic component of the Lorentzian shape.
We fit all the parameters together (slope and normalization of the power law, normalization and the broadening of the Fe II template, the position, normalization and FWHM of the Mg II) since in case of such a short data range we do not have any possibility to find parts of the spectrum unaffected by Fe II and fit first the underlying power law.
The narrow spectral range makes the fits not strongly constraining the Fe II templates but the fits are very satisfactory. The results of the spectral fitting are shown in Figure 13. The best solution corresponds to , shown in the middle panel of Figure 13. The best fit FWHM of the Fe II (5400 km s) is broader than the FWHM of Mg II (4280 km s) but it comes with a large error. Higher turbulent velocity (100 km s) gives fits with somewhat higher reduced , and pushes the FWHM of Fe II toward even higher values, at the same time increasing the Fe II intensity. This results from considerable degeneracy between the Fe II normalization and the underlying power law. A lower value of the turbulent velocity (20 km s) gives narrower Fe II lines but the quality of the fit is much worse.
For comparison, we show two other fits, for higher ( (cm s) = 19) and lower ( (cm s) = 18) ionization fluxes in the upper and lower panels of Figure 13, respectively. The fit quality is good in all these cases ( differences ), and the value of the EW of the Mg II line is similar (31.2 Å, 30.2 Å and 28.4 Å), but slowly rising towards larger ionization fluxes. We calculated also EW(Fe II) but only in the fitted range. EW(Fe II) is more affected by the change of the ionizing flux (the corresponding values are 37.9 Å, 31.2 Å and 15.9 Å), and all these values are larger than achieved with the use of different templates by Zajaček et al. (2023b). Such a large change is related to a corresponding change in the slope and normalization of the underlying power law.
The chosen fit density ( (cm) = 12) is consistent with the best fit theoretical template d12-m20-20-5.dat selected from the collection of Bruhweiler & Verner (2008) in our previous papers, including Prince et al. (2023). However, the ionization flux is now lower by more than two orders of magnitude.
We compared the requested value for the ionizing flux from Fe II modelling with the expected ionizing flux based on SED of this object and the Mg II time delay of 302.6 days for Mg II (Zajaček et al., 2020) but the obtained ionization flux from the SED best fit presented in Figure 9 of Zajaček et al. (2020) gave only , since the spectrum appeared to decrease rapidly in the UV band.
The narrow available spectral range on one hand guarantees that a range of templates fits well the spectrum but it opens a question of how reliable is such a decomposition.
3.8.2 SDSS RM102
As a second example, we selected the quasar SDSS J141352.99+523444.2 (SDSS RM102) from the SDSS Reverberation Mapping (SDSS-RM) catalogue (Shen et al., 2019) considering a high signal-to-noise, moderate Fe II emitter (), and a reliable time lag for H or Mg II (maximum Pearson correlation coefficient ). The spectrum of the source covers a broad wavelength range (2000-5500 Å). This allowed us to cover both the rest-frame optical Fe II as well as UV Fe II and to test the feasibility of using the same Fe II template to fit the entire UV-to-optical spectral range. The size of its BLR region has been estimated by different methods, in this paper we took the one estimated by the Interpolated Cross-Correlation Function (ICCF), and days in the observed-frame (Shen et al., 2023). Using the SDSS Science Archive Server (SAS) from the DR16333https://dr16.sdss.org/optical/spectrum/search, we found 70 spectroscopic epochs observed during four years. All the spectra were combined in a single one to obtain a high signal-to-noise spectrum, S/N=30 and 35 at 3000Å and 4000Å, respectively. Due to the presence of spikes around 5100Å, it was not possible to estimate the S/N at this wavelength.
In the spectral fitting, we considered all the emission lines reported by Vanden Berk et al. (2001) between 2000 and 5500 Å, 31 emission lines in total. We also included a model of the Balmer Continuum with an optical depth of 1 plus the High-Order Balmer lines (Bernal et al in prep.). The continuum was fitted with a single power law. To decrease the number of free parameters, we fixed the same FWHM for each family of emission lines: Balmer lines, Helium transitions, forbidden lines (narrow emission lines), and coronal lines. Flux intensities, FWHM, and shift sum a total of 63 free parameters. We applied Gaussian smearing to the UV and optical Fe II template and a single flux intensity value was taken for the UV and optical wavelengths. In the first attempt, we noticed that it was not possible to fit the UV and optical Fe II contribution with the same flux intensity. Thus, we performed a careful fitting of all the other spectral features and what remained can be considered as the Fe II component. Due to the overlapping of the emission lines and the Fe II pseudocontinuum in some ranges, the flux intensity of the emission line would be underestimated. So, we considered the semiempirical UV/optical Fe II templates implemented in the fitting code PyQSOFit (Guo et al., 2018) to constrain the intensity of the emission lines. As we previously noticed, even with these empirical Fe II templates, a different flux intensity value is needed to fit the Fe II contribution. The flux intensity factor has a difference of 3.5 between the UV and the optical Fe II contribution. The final spectral fitting is shown in Figure 14. The potential Fe II contribution is shown at a zero level. As a reference, we also plotted the UV (dotted line) and optical (dashed-line) Fe II templates implemented in PyQSOFit. There is good agreement in the optical Fe II, while slight variations are seen in the UV part.
The new spectral fitting indicates a =1.24, which is approximately twice the value reported by Shen et al. (2019). On the other hand, this source shows a moderate Eddington ratio =0.329 (Shen et al., 2019). The weak contribution of the doublet [OIII] suggests that RM102 has a moderate Eddington ratio and is close to the xA sources (Marziani, 2023).
To locate the source in the plane, we measured three ratios: red to blue wings of UV Fe II, red to blue wings of optical Fe II, and UV to optical Fe II. The first two ratios were measured in the same wavelength range as used to create the general maps in Figure 2 and Figure 5. The last ratio had to be modified since the data cover the wavelength range only up to 2000 Å. The measured ratios are given in Table 4. Looking at the plots in Section 3 above we could not easily identify the most likely parameters of the successful template. We thus created a new map, shown in Figure 15, where all the three ratios given in Table 4 are shown as contour plots, each for all 6 turbulent velocities.
wavelength range 1 | wavelength range 2 | ratio 1 to 2 |
2800-3090 Å | 2650-2800 Å | 1.939 |
5150-5350 Å | 4434-4684 Å | 0.713 |
2000-3090 Å | 4000-5350 Å | 2.151 |
We see that the three families of lines in Figure 15 never cross. This means that there is no single template in our collection which could fit the data for the quasar RM102. The blue lines representing the ratio of red to blue wings of optical Fe II occupy a narrow range of high ionization flux values. This reflects the fact that the measured ratio is low and the corresponding solutions are located in the upper part of the general diagram (see Figure 5). On the other hand, the UV wings ratio in general forms a two-family solution for low values (see Figure 2), but for high values as the one measured in our source there is only an island located at low ionization flux and low density. The global UV to optical ratio (although measured in a slightly different wavelength range than the generic Figure 1) also points toward low density and low ionization flux. The results for densities of the order of cm and moderate ionization fluxes of the order of cm s are similar to the parameter range concluded for HE 0413-4031, but the optical wing ratio remains unexplained.
We estimated the ionization flux from this source at the distance corresponding to the BLR position as given by the H line delay. However, the SDSS spectrum does not cover the spectral range above 1 Rydberg, needed for this purpose. We thus extrapolated the spectrum of RM102 using the AGN SED template from CLOUDY. The observed spectrum is only marginally bluer in the optical band. Matching the two spectra at UV part we obtained the ionization flux of cm s. We also tested another way of extension, using the HST composite spectrum of Zheng et al. (1997) which covers the rest frame range from 350 Å up to 2200 Å, and they give the broken power law fit. Matching this broken power law to the shortest wavelengths with our spectrum we obtained the ionization flux of cm s. The two values are comparable, they eventually favour the value implied by the red-to-blue optical Fe II ratio, although at (rather likely) higher densities than cm they are not in full agreement with any of the ratios. However, some composites predict a stronger decrease of the quasar far-UV flux towards shorter wavelengths (see Cai & Wang, 2023, and the references there).
Overall, Figure 15 implies that a single Fe II template is unlikely to fit well the whole spectrum. In particular, the mismatch in the values of the ionization flux is by a factor of 100. The metallicity is not affecting the ratios strongly enough. We thus set the density at the likely value of cm, and we took two approaches to set the ionizing flux. In the first case, we allowed for a discontinuity in the ionizing flux at 4000 Å, where the Fe II emission is very low, and we used low ionizing flux in UV band () and high ionizing flux () in the optical band. Separate fitting of the Fe II and optical is usually done when modelling a specific source. The result of this approach is shown in Figure 16. In the second approach, we allowed for a linear change of the ionizing flux with the wavelength in the log-log scale. This result is shown in the same plot (magenta line). None of these attempts are fully satisfactory. However, the discontinuous model seems to work much better. Still, in both cases, we face a considerable mismatch between the model and the data in the region 3000 - 3500 Å. Semi-empirical models do include considerable emission there (see Figure 14, dotted line). The spike close to 2000 Å is most likely a Fe III emission. We did not try a more complex approach in the form of the Locally Optimized Cloud (LOC) model of Baldwin et al. (1995) since we do not include Fe III emission which may be important for exact data fitting. We address this issue and additional likely effects in the discussion.
4 Discussion
We used the most recent version of the code CLOUDY C23.00 to construct the grid of theoretical Fe II templates covering the broad UV to NIR range from 1000 Å to 10000 Å. The templates represent single-zone emitters, parametrized by the local density and ionization flux. We consider a range of turbulent velocities and for a fixed turbulent velocity of 20 km we calculated the results for different metallicities. For the other turbulent velocity the considered metallicity is solar.
The advantage of the theoretical templates is that they predict the Fe II emission underlying the strong emission lines while purely observational templates, modelling first the lines, force low Fe II emission at the position of lines like H or Mg II. In addition, the construction of the observational templates is thus most successful for the objects with extremely narrow emission lines like I Zw 1 (Vestergaard & Wilkes, 2001) or Mrk 493 (Park et al., 2022). Line widths of these spectra imply high Eddington ratio, strong Fe II emission and the incident SED shape not necessarily representative for the majority of quasars.
The availability of the Fe II templates in the full spectral range will address better the issue of the production of the Fe II in the BLR. In the case of the observational data covering a narrow spectral range, the templates allow us to fit the data and obtain an estimate of the density and the ionization flux in the BLR, as we showed in Section 3.8.1. However, our attempt to use these templates for a quasar spectrum covering both optical and UV ranges was not successful (see Section 3.8.2). The problem seems generic since the selected quasar is rather a typical object, and the issue of the UV to optical ratio in the models and the data has been present in the literature for many years.
Already the early papers asked the question of whether the Fe II emission actually forms as a result of the irradiation by the central source or due to mechanical heating. For example, Joly (1981) in her paper argued that either mechanical heating is the dominating channel of production, or the clouds must be very optically thick, in which case the irradiation flux is thermalized and the mechanism is similar to the mechanical heating. However, those early models were not suitable for column densities larger than cm.
Verner et al. (2003) calculated the predicted ratio of the UV to optical Fe II emission from their model of the Fe II atomic transitions. The adopted ranges were 2000–3000 Å and 4000–6000 Å range, the range is comparable although not identical to the one studied in Section 3.8.2. The UV/optical ratio obtained in this paper is similar to what we measure from RM102 in the case of collisional heating, but radiative heating gives much higher values, 8 - 11, depending on the number of transitions. These values were obtained under a turbulent velocity of 1 km s. They noted that Fe II optical emission flux is more sensitive to abundance than the Fe II (UV) band. Conversely, the Fe II (UV) band is more sensitive to microturbulence than the Fe II optical band.
Baldwin et al. (2004) investigated the Fe II emission in the UV (2200-2800 Å) and optical (4450-4750 and 5080-5460 Å) ranges. They obtained the observed Fe II UV/optical ratio of 1-2, which is significantly lower than the model-predicted ratio of 40-50. Using the LOC model, they found that the observed UV Fe II emission requires a high microturbulence velocity ( 100 km/s). They examined collisional heating as an alternate excitation mechanism, however only the Fe II lines were visible in the UV range, which has the disadvantage of adding another specialised region to the already complicated structure of the inner region of AGN. They noticed that other complex ions, such as Ni II, might have created similar patterns despite their lower abundance, but their atomic data was not available. As a result, they favoured a microturbulence-based model for explaining UV Fe II emissions and concluded that the optical Fe II emission originates from an entirely different region compared to the UV Fe II emission.
The problem of matching the theoretical templates over the full optical to UV spectral range is not likely due to the use of a single-zone approximation, particularly as some of the papers mentioned above used a multi-zone approach. The standard LOC model can in principle represent both the broad range of radii as well as the broad range of densities but may not be actually appropriate to solve the physical problem.
The radially very extended emission is unlikely. Kovačević-Dojčinović & Popović (2015) argued that there is a kinematic correlation between the UV and the optical Fe II emissions, although they reported considerable positive shift of the UV Fe II with respect to the systemic redshift for the majority of their sources (their Fig. 10). The radius-luminosity relation discussed by Zajaček et al. (2023b) for both spectral ranges looks roughly similar, so the two regions cannot be considerably shifted. Furthermore, the possibility of measuring the Fe II delay in the optical and UV bands contradicts the major collisional origin of the emission, but some contributions cannot be ruled out. The atomic content of the recent version C23.00 of CLOUDY used in this paper does not seem questionable, and apparently, the systematic improvement in the data did not solve the long-standing issue of UV and optical discrepancy. The metallicity dependence also does not seem to be strong enough to account for the issue.
There are thus two remaining issues which can solve the problem but they were not modelled in this paper since our primary goal was to provide a set of theoretical templates for further use.
The first weakness in the simulation is the use of a constant density assumption. Such an approach is widely employed, yet it is not physically warranted. The higher temperature of the illuminated face of the cloud and the considerably lower temperature of the dark side would create an enormous pressure gradient which could disrupt the cloud during the dynamical timescale. Instead, the illuminated cloud must be in pressure equilibrium which also takes into account the radiation pressure (Różańska et al., 2006; Baskin et al., 2014). Overall, the bright face has a much lower density, and the density is rising towards the dark side. Clouds in pressure equilibrium were considered in several papers. Some of these computations were done in the context of the warm absorber (Różańska et al., 2006; Adhikari et al., 2015, 2019) but this medium may be the same as BLR. The change of the density and temperature is not like a power law but rather like a rapid transition (see e.g. Fig. 1 of Baskin et al. 2014, or Fig. 3 of Adhikari et al. 2019), so the LOC model would not reproduce it well, and eventually, a discontinuous jump in density and ionization parameter might be a better approximation. Computations of clouds under constant pressure are possible within CLOUDY environment, and this could be possibly done in the future.
The second weakness is the assumption of isotropic emission. If the clouds are few and randomly distributed around the central source, then we will see clouds at all angles with respect to their orientation towards the source. The entire contribution summed up is well represented by the isotropic emission. If cloud distribution is flattened, but our line of sight is not passing through the cloud region the dark sides may contribute a little stronger to the total emission in terms of the effective surface but most of the cloud emission comes from the bright sides. Such expectations were given, for example, by Czerny & Hryniewicz (2011) in the context of the time delay. However, the bright sides contribute more to the isotropic part, therefore overexposure of the cloud’s bright sides may not have a significant impact on Fe II emission. However, Ferland et al. (2009) suggested that we predominantly see the dark side of the clouds. They connected this to the Fe II emission in the optical band showing predominantly the redshift (Hu et al., 2008) and thus implying the inflow. Also in Panda et al. (2018), in Section 4.3, they discussed the issue that the optical Fe II is produced predominantly in the dark side of a cloud, and the dominance of the dark side helped to reproduce ratio without metallicity enhancement. However, the requested geometrical setup is not simple since the line of sight does not cross the BLR. It may require a highly ionized absorber, between the black hole and the BLR, almost ‘invisibly’ shielding the well-irradiated bright sides of clouds.
5 Summary
One of the most prominent characteristic features of the quasar spectrum is the iron emission. In this work, we examine the properties of Fe II emission using CLOUDY version C23.00 and generate new theoretical Fe II templates that can be used to model Fe II emission in the quasar spectra covering the wavelength range 1000-10000 Å. We study the impact of microturbulence and metallicity on iron emission and compare our model predictions with observational data. The primary outcomes of our study are outlined below.
1.
The long-standing problem of a mismatch between the expected and observed values of Fe II UV to optical ratio under the single zone approximation could not be resolved by the updated atomic database available in CLOUDY C23.00.
2.
Microturbulence has less impact on the UV to optical Fe II emission ratio.
3.
An additional apparent velocity shift of up to 1000 km/s can be added to the spectrum by microturbulence.
4.
The ionisation flux has a more significant impact on the ratio of red-to-blue parts of optical Fe II around H than does the gas density.
5.
High-density BLR ( (cm) = 12 - 13) and extremely high radiation fluxes are needed for the highly accreting (xA) sources.
6.
The ratios of Fe II emissions are not significantly affected by the metallicity.
7.
The observed UV-to-optical spectra of a quasar cannot be well fitted by a single theoretical Fe II template generated assuming fixed column density and isotropic cloud emission.